Two tanks, X and Y, are filled to capacity with jet fuel. Tank X holds 600 gallons more than tank Y. If 100 gallons were to be pumped out from each tank, tank X would then contain 3 times as much fuel as tank Y. What is the total number of gallons of fuel in the two full tanks?
Answer by richwmiller(17219)

x=y+600,
x-100=3*(y-100)
y+600-100=3*(y-100)
y+500=3y-300
800=2y
y=400
x=1000
check
900=3*300
ok